# LLM4AD: Large Language Models for Autonomous Driving — Concept, Review, Benchmark, Experiments, and Future Trends

### v4

@@ III-A1 Supervised Fine-Tuning
Supervised fine-tuning (SFT) is a well-established technique in machine learning that adapts a pre-trained model to a downstream task by further training it on task-specific data. In our context, SFT allows LLMs to leverage their world knowledge acquired during pre-training while specializing it for driving tasks. However, full-weight SFT is a highly resource-intensive process, especially for LLMs, requiring substantial computational power and memory to store and update the model’s billions of parameters.
@@ III-B2 Output
In Tab. I, we use the term signal to refer to control signals represented as numerical values. For instance, DriveGPT-4 predicts control signals, including the ego vehicle’s speed and turning angle, for the subsequent step.
@@ IV-A1 Dataset
The LaMPilot-Bench dataset consists of 4.9K semi-human-annotated traffic scenes, with a subset of 500 samples split as the test set. Each data sample includes a high-level task description (instruction), an initial state for initializing the simulator, and goal state criteria aligned with the instruction. The dataset covers diverse driving scenarios, as shown in Tab. III. Our “semi-human-annotated” protocol involved a two-stage process. First, for each scenario, LaMPilot-Bench includes various situations, such as different ego vehicle initial positions and states, specific tasks (turning left/right or going straight), the number of other vehicles, and their positions and states. The driving model parameters for other vehicles are randomly initialized, and each scenario is assigned a random seed. Second, these scenarios were presented to a team of human annotators. These annotators were tasked with writing a clear, natural language instruction (e.g., “Move to the middle lane and drive at 30 mph”) that corresponded to the programmatic goal. To ensure the quality and consistency of these annotations, we then implemented a quality control pass. A separate team of senior evaluators reviewed the annotations, checking for ambiguity, grammatical errors, and, most importantly, a clear alignment between the natural language instruction and the scenario’s programmatic goal. Any instruction that was found to be ambiguous or did not uniquely map to the intended goal was discarded and sent back for re-annotation.
The LaMPilot-Bench simulator is built upon HighwayEnv [leurent_environment_2018], a widely used platform for autonomous driving research and tactical decision-making, which offers various driving models and simulates realistic multi-vehicle interactions. We extend HighwayEnv with interfaces suitable for LLM-based agents and implement custom intersections to diversify the driving scenarios.
@@ IV-A2 Evaluator
Figure 3: Distribution of instructions in the LaMPilot-Bench dataset.
@@ IV-A3 Baselines and Results
We establish several baselines to evaluate the performance of LLM-based agents across the benchmarks, as presented in Tab. IV for LaMPilot-Bench.
@@ IV-B1 Dataset
In addition to LaMPilot-Bench in HighwayEnv, we also utilize the CARLA Leaderboard 1.0 benchmark [carla_team_autonomous_2020, dosovitskiy_carla_2017] to evaluate LLM-based agents in a more realistic and complex driving environment. The Leaderboard consists of 76 routes (50 for training, 26 for testing), totaling over 170 km across six towns. It also includes predefined scenarios covering ten types of challenging traffic situations, such as control loss and traffic negotiation. To test the instruction-following capabilities of LLM-powered systems, we modify the Leaderboard setup to provide natural language navigation instructions based on the agent’s current position instead of GPS coordinates. To eliminate perception influence on our results, we craft the driving context information using a language generator that leverages CARLA’s privileged information. Executing an LMP in CARLA involves using Python’s exec function, which takes the LMP code as an input string and two dictionaries defining the execution scope: (i) apis, containing all driving APIs the code may call, and (ii) local_vars, initially empty but later holding a generator variable named policy once executed. The Leaderboard evaluates agent performance using three key metrics:
@@ IV-B2 Baselines and Results
Figure 4: An overview of NuPlanQA. NuPlanQA comprises nine subtasks across three skill areas to support the context-aware analysis of traffic scenes. The proposed baseline, BEV-LLM, is trained on NuPlanQA-1M using historical frames, BEV features from multi-view images, and control signals as inputs. Finally, we evaluate MLLMs using NuPlanQA-Eval, a multiple-choice QA benchmark for driving scene understanding.
We compare the human-feedback baseline with reinforcement learning experts that have access to privileged information, such as Roach [zhang_end--end_2021] and TCP [wu_trajectory-guided_2022] in the CARLA leaderboard [ma_learning_2024], as shown in Tab. V. The comparison aims to provide a reference to previous works. Based on these results, we observe the following:
@@ IV-C1 Dataset
While LaMPilot-Bench and CARLA Leaderboard evaluate LLM-based agents in closed-loop simulation, they are limited to low-level control policies and rule-based task execution. To further assess the high-level reasoning and perception capabilities of MLLMs in realistic driving environments, we incorporate the recently proposed NuPlanQA benchmark [park2025nuplanqalargescaledatasetbenchmark].
@@ IV-C2 NuplanQA Baseline and Results
We extract features from multiple layers of $\hat{V}_{\text{BEV-att}}$ (every fourth layer and the final hidden state [grattafiori2024llama3herdmodels]) and concatenate them to create a robust multi-scale representation. A multi-modal projector then aligns this representation with the language model’s text embedding space.
@@ IV-D Summary
The LaMPilot-Bench dataset and the CARLA Leaderboard 1.0 benchmark provide a comprehensive evaluation framework for assessing the instruction-following capabilities and performance of LLM-based agents in autonomous driving. The experiments highlight the potential of LLMs in enhancing the decision-making capabilities of autonomous vehicles when combined with human feedback and domain-specific APIs. As LLMs continue to evolve and integrate with more modalities, we envision their significant contribution to the development of safe, efficient, and user-friendly autonomous vehicles. Future research should focus on addressing the challenges of real-time decision-making, reducing collision rates, and incorporating multi-modal inputs to further enhance the performance and robustness of LLM-based autonomous driving systems.
@@ V Real-Vehicle Experiments: On-Cloud LLM for Personalized Decision-Making
Figure 6: The proposed framework for personalized decision-making with an on-cloud large language model in real-vehicle experiments.
To further assess the effectiveness of the LLM4AD system and validate its applicability in real-world scenarios, we integrate LLMs into practical autonomous driving systems, introducing a framework called Talk2Drive. Similar to the concept in Sec. II, our system has outstanding command understanding and reasoning performance. It also can provide a personalized driving experience by leveraging records of previous interactions. The demo for our real-vehicle experiment is publicly available [^1] Real-vehicle experiment demo video with on-cloud LLM: https://youtu.be/4BWsfPaq1Ro.
@@ V-A1 Problem Statement
In Talk2Drive, we follow the concept we proposed in Sec. II. The model aims to translate verbal commands into executable control sequences for the vehicle. The cloud-based LLM acts as a translating function $f:I\rightarrow P$ that generates corresponding LMPs as the policy ($P$) for maneuvers.
For personalization, we need human feedback on execution $F(I,P)$ to evaluate if the generated policy addresses the preferences of the driver or the passengers. The extra memory module (see V-A5), which includes instructions, LMPs, and their evaluations, allows the LLM to learn user preferences. Therefore, there are two stages in the Talk2Drive workflow:
@@ V-A2 Command Translation and Contextual Data Integration
Simultaneously, LLMs access additional cloud-based real-time environment data, including weather updates, traffic conditions, and local traffic rules. In our product, LLMs can be informed by the weather information through Openweather API [openweather], the map information (such as road type and speed limits) through OpenStreetMap API [OpenStreetMap], and traffic information through TomTom API [tomtom]. We use a predefined structured language generator for both contextual data ($C$) and system messages ($S$), and our system will supply the appropriate values (the red contents) to the generator based on the context information, as shown below:
@@ V-A3 Processing and Reasoning with LLMs
Once a command is translated into texts, it is uploaded to LLMs hosted on the cloud. In our experiments, we utilize GPT-4 [openai_gpt-4_2023] as our LLM and access it through ChatGPT API. The LLMs engage in a reasoning process to interpret the instruction ($I$), the contextual data ($C$), the system messages ($S$), and the history data ($H$). A simplified example of $S$ is presented below:
@@ V-A4 Actionable Code Generation
One simple example codes are shown as follows:
@@
In the meantime, LLMs also generate thoughts $R$ to explain why they make a particular decision. According to  [wei_chain--thought_2022], even though the thoughts $R$ are not executed in our systems, they help LLMs make more reasonable and feasible decisions. An example of the thought $R$ is shown below:
@@
The generated LMPs ($P$) are then sent back from the cloud to the vehicle’s electronic control unit (ECU), where they are executed. Additionally, we also set two kinds of safety checks for the generated LMPs. The first check involves verifying the format of the generated code. If the code does not conform to a valid format, our system will not execute or respond to the code. The second safety check is parameter verification, which assesses the feasibility and safety of the parameters within the context of the given policy, preventing the execution of potentially risky policies. For example, if the generated code instructs our agent to move to the left lane while the ego agent is already in the leftmost lane, our system will identify this abnormality and prevent the execution.
@@ V-A5 Memory Module and Personalization
Given the adaptive nature of LLMs, if users respond differently to similar commands, the LLMs will prioritize the most recent response as a reference point for their current decision-making process. When a command from the users is issued, the LLMs access the memory module and take stored information ($H$) as part of the input prompts for the decision-making process. Additionally, each user has their own profile in the memory module, ensuring that our framework can deliver personalized driving strategies for everyone.
@@ V-B Autonomous Driving System
As shown in Fig. 8, we use an autonomous vehicle to conduct real-world experiments, which is a drive-by-wire enabled 2019 Lexus RX450h with perception sensors, localization module, and communication module. We deploy the open-source autonomous driving software Auoware.AI [autoware] with ROS Melodic in Ubuntu 18.04. We employ 3D-NDT [3Dndt] for mapping and localization, and we utilize pure pursuit [pure_pursuit] for trajectory following. Network communication is enabled by the Cradlepoint IBR900 Series Router, which is equipped with an AT&T sim card with a 4G-LTE cellular connection.
@@ V-C1 Scenarios Setup and Participants
Our experimental trials include three different scenarios: highway, intersection, and parking lot.  [^2] The experiments conducted in this study satisfy all local traffic guidelines and guarantee the safety of the participants. A human always sits in the driver’s seat of the autonomous vehicle to monitor its status and get ready to take over. The field experiments for the highway and intersection scenarios are conducted at a proving ground in Columbus, IN, USA, to validate the efficacy of the proposed Talk2Drive framework. The highway scenario involves a three-lane highway, while the intersection contains a two-way junction. Additionally, the parking lot scenario is evaluated at the North Stadium Parking Lot in West Lafayette, IN, USA. The overview visualization and statistics of the test tracks are shown in Fig. 9, while the visualization for the experiments is shown in Fig. 10. In our study, we recruit a diverse range of drivers, including seven individuals of different genders (61.4% male, 28.6% female), ages (mean=26.71, std=4.11), and driving experiences(mean=6.79, std=5.08). All participants have valid driving licenses.
@@ V-C2 Input Instructions
Given that conventionally indirect strategies primarily modify the politeness level of command of direct strategies, we instead divide the non-conventionally indirect requests into two categories based on the strength of the hints. We define our levels of directness in Tab. VII as our way to classify varying degrees of implicitness in the commands. To gather data, we request our test drivers to generate commands based on their normal speaking preferences and subsequently categorize their commands into our defined levels. Examples of the commands generated are also presented in Tab. VII.
@@
Driving
Secnario
Expected Driving
Behavior
Command
Directness
@@ V-C3 Evaluation Metrics
where $\tau_{c}$ is the critical threshold. While 2.0 s is a common simulation margin, a 1.5 s threshold was selected for this real-world test based on existing research results on TTC value [van1994time]. For other metrics like speed variance, acceleration, and jerk, the thresholds are related to individual perception and would vary in different driving scenarios. Therefore, we define the score to quantify the corresponding performance, with a higher score indicating better performance relative to the baseline value. For example, the sub-score for mean absolute jerk $|\bar{J}|$ can be denoted as:
@@
where $\gamma$ is the sensitivity factor which is set empirically. The sub-scores for mean absolute acceleration and speed variance are obtained in the same way.
LLMs’ Latency for Time Efficiency: Latency $Lcy$ measures the response time of the LLM, which is important for time-sensitive vehicle applications. To measure latency, we calculate the time difference between the moment a command is sent to the LLMs and when the LLMs return the LMPs.
@@ V-D1 The Validation of Driving Performance of Talk2Drive
To validate the driving performance of the Talk2Drive system, we conduct comprehensive experiments to compare the driving performances between the baseline and the Talk2Drive through the metrics defined in Sec. V-C3. As discussed earlier, the baseline values are from the average data of human drivers. The average driving performance for each scenario within each command directness is shown in Tab. VIII. Note that for each expected driving behavior, the driving pattern is distinct. Therefore, a baseline is created for every driving behavior separately.
The driving score is the weighted sum of all the safety and comfort metrics, reflecting the overall driving performance. The driving scores are calculated and recorded in Tab. VIII, with green texts highlighting scores higher than the baseline of the same driving behavior and red texts showing the lower scores. As shown in Tab. VIII, regardless of the command’s directness or the human’s behavior, nearly all driving scores are better than the baseline, showing that our framework consistently provides relatively safe and comfortable driving experiences for humans in all scenarios.
@@
The latency of the internet remains mostly consistent (around 300 ms) under the same network conditions. The empirical results from our tests reveal that the latency for the most responsive LLM falls between approximately 1.2 to 1.8 seconds. While the observed 1.2 s to 1.8 s latency is acceptable for high-level, non-urgent decisions (e.g., responding to “drive more conservatively”), we must emphasize that this latency makes the cloud-based LLM entirely unsuitable for any safety-critical, time-sensitive maneuvers (e.g., emergency braking or collision avoidance.) This finding suggests considerable potential for integration into certain real-time applications where such response times are within acceptable limits. Specifically, LLMs demonstrate strong suitability for making high-level decisions or modifying driving behaviors in response to human commands. These scenarios require timely responses but do not typically demand extremely low latency.

### v5

@@ III-A1 Supervised Fine-Tuning
Supervised fine-tuning (SFT) is a well-established technique in machine learning that adapts a pre-trained model to a downstream task by further training it on task-specific data. In this paper’s context, SFT allows LLMs to leverage their world knowledge acquired during pre-training while specializing it for driving tasks. However, full-weight SFT is a highly resource-intensive process, especially for LLMs, requiring substantial computational power and memory to store and update the model’s billions of parameters.
@@ III-B2 Output
In Tab. I, the term signal is used to refer to control signals represented as numerical values. For instance, DriveGPT-4 predicts control signals, including the ego vehicle’s speed and turning angle, for the subsequent step.
@@ IV-A1 Dataset
The LaMPilot-Bench dataset consists of 4.9K semi-human-annotated traffic scenes, with a subset of 500 samples split as the test set. Each data sample includes a high-level task description (instruction), an initial state for initializing the simulator, and goal state criteria aligned with the instruction. The dataset covers diverse driving scenarios, as shown in Tab. III. Our “semi-human-annotated” protocol involved a two-stage process. First, for each scenario, LaMPilot-Bench includes various situations, such as different ego vehicle initial positions and states, specific tasks (turning left/right or going straight), the number of other vehicles, and their positions and states. The driving model parameters for other vehicles are randomly initialized, and each scenario is assigned a random seed. Second, these scenarios were presented to a team of human annotators. These annotators were tasked with writing a clear, natural language instruction (e.g., “Move to the middle lane and drive at 30 mph”) that corresponded to the programmatic goal. To ensure the quality and consistency of these annotations, a quality control pass is then implemented. A separate team of senior evaluators reviewed the annotations, checking for ambiguity, grammatical errors, and, most importantly, a clear alignment between the natural language instruction and the scenario’s programmatic goal. Any instruction that was found to be ambiguous or did not uniquely map to the intended goal was discarded and sent back for re-annotation.
The LaMPilot-Bench simulator is built upon HighwayEnv [leurent_environment_2018], a widely used platform for autonomous driving research and tactical decision-making, which offers various driving models and simulates realistic multi-vehicle interactions. the HighwayEnv is extended with interfaces suitable for LLM-based agents and implement custom intersections to diversify the driving scenarios.
@@ IV-A3 Baselines and Results
Several baselines are estabilished to evaluate the performance of LLM-based agents across the benchmarks, as presented in Tab. IV for LaMPilot-Bench.
@@ IV-B1 Dataset
In addition to LaMPilot-Bench in HighwayEnv, the CARLA Leaderboard 1.0 benchmark [carla_team_autonomous_2020, dosovitskiy_carla_2017] is used to evaluate LLM-based agents in a more realistic and complex driving environment. The Leaderboard consists of 76 routes (50 for training, 26 for testing), totaling over 170 km across six towns. It also includes predefined scenarios covering ten types of challenging traffic situations, such as control loss and traffic negotiation. To test the instruction-following capabilities of LLM-powered systems, the Leaderboard setup is modified accordingly to provide natural language navigation instructions based on the agent’s current position instead of GPS coordinates. To eliminate perception influence on our results, a language generator is used to craft the driving context information that leverages CARLA’s privileged information. Executing an LMP in CARLA involves using Python’s exec function, which takes the LMP code as an input string and two dictionaries defining the execution scope: (i) apis, containing all driving APIs the code may call, and (ii) local_vars, initially empty but later holding a generator variable named policy once executed. The Leaderboard evaluates agent performance using three key metrics:
@@ IV-B2 Baselines and Results
Figure 2: An overview of NuPlanQA. NuPlanQA comprises nine subtasks across three skill areas to support the context-aware analysis of traffic scenes. The proposed baseline, BEV-LLM, is trained on NuPlanQA-1M using historical frames, BEV features from multi-view images, and control signals as inputs. Finally, MLLMs are evaluated using NuPlanQA-Eval, a multiple-choice QA benchmark for driving scene understanding.
The human-feedback baseline with reinforcement learning experts that have access to privileged information, such as Roach [zhang_end--end_2021] and TCP [wu_trajectory-guided_2022] in the CARLA leaderboard [ma_learning_2024] are compared, as shown in Tab. V. The comparison aims to provide a reference to previous works. Based on these results, the follow conclusions are observed:
@@ IV-C1 Dataset
While LaMPilot-Bench and CARLA Leaderboard evaluate LLM-based agents in closed-loop simulation, they are limited to low-level control policies and rule-based task execution. To further assess the high-level reasoning and perception capabilities of MLLMs in realistic driving environments, the recently proposed NuPlanQA benchmark is incorporated [park2025nuplanqalargescaledatasetbenchmark].
@@ IV-C2 NuplanQA Baseline and Results
The features from multiple layers of $\hat{V}_{\text{BEV-att}}$ (every fourth layer and the final hidden state [grattafiori2024llama3herdmodels]) are extracted and concatenated to create a robust multi-scale representation. A multi-modal projector then aligns this representation with the language model’s text embedding space.
@@ IV-D Summary
The LaMPilot-Bench dataset and the CARLA Leaderboard 1.0 benchmark provide a comprehensive evaluation framework for assessing the instruction-following capabilities and performance of LLM-based agents in autonomous driving. The experiments highlight the potential of LLMs in enhancing the decision-making capabilities of autonomous vehicles when combined with human feedback and domain-specific APIs. As LLMs continue to evolve and integrate with more modalities, their significant contribution is envisioned to the development of safe, efficient, and user-friendly autonomous vehicles. Future research should focus on addressing the challenges of real-time decision-making, reducing collision rates, and incorporating multi-modal inputs to further enhance the performance and robustness of LLM-based autonomous driving systems.
@@ V Real-Vehicle Experiments: On-Cloud LLM for Personalized Decision-Making
Figure 4: The proposed framework for personalized decision-making with an on-cloud large language model in real-vehicle experiments. On the cloud side, human verbal commands are transcribed and combined with real-time environmental context—such as weather, traffic rules, and road conditions—to serve as inputs for a large language model. This model also accesses historical interaction data to generate executable code policies and reasoning thoughts tailored to the user. These outputs are transmitted to the physical vehicle, where an electronic control unit executes them within the autonomous stack to adjust trajectory generation and motion control parameters. The process concludes with a feedback loop, where human evaluations of the driving maneuvers are stored in a memory module to continuously refine the system’s personalization capabilities.
To further assess the effectiveness of the LLM4AD system and validate its applicability in real-world scenarios, we integrate LLMs into practical autonomous driving systems, introducing a framework called Talk2Drive. Similar to the concept in Sec. II, this system has outstanding command understanding and reasoning performance. It also can also provide a personalized driving experience by leveraging records of previous interactions. The demo for the real-vehicle experiment is publicly available [^1] Real-vehicle experiment demo video with on-cloud LLM: https://youtu.be/4BWsfPaq1Ro.
@@ V-A1 Problem Statement
In Talk2Drive, followed by the concept proposed in Sec. II. The model aims to translate verbal commands into executable control sequences for the vehicle. The cloud-based LLM acts as a translating function $f:I\rightarrow P$ that generates corresponding LMPs as the policy ($P$) for maneuvers.
For personalization, the human feedback on execution $F(I,P)$ is needed to evaluate if the generated policy addresses the preferences of the driver or the passengers. The extra memory module (see V-A5), which includes instructions, LMPs, and their evaluations, allows the LLM to learn user preferences. Therefore, there are two stages in the Talk2Drive workflow:
@@ V-A2 Command Translation and Contextual Data Integration
Simultaneously, LLMs access additional cloud-based real-time environment data, including weather updates, traffic conditions, and local traffic rules. In this product, LLMs can be informed by the weather information through Openweather API [openweather], the map information (such as road type and speed limits) through OpenStreetMap API [OpenStreetMap], and traffic information through TomTom API [tomtom]. We use a predefined structured language generator for both contextual data ($C$) and system messages ($S$), and this system will supply the appropriate values (the red contents) to the generator based on the context information, as shown below:
@@ V-A3 Processing and Reasoning with LLMs
Once a command is translated into texts, it is uploaded to LLMs hosted on the cloud. In experiments, GPT-4 [openai_gpt-4_2023] is utilized and access it through ChatGPT API. The LLMs engage in a reasoning process to interpret the instruction ($I$), the contextual data ($C$), the system messages ($S$), and the history data ($H$). A simplified example of $S$ is presented below:
@@ V-A4 Actionable Code Generation
One simple example code is shown as follows:
@@
In the meantime, LLMs also generate thoughts $R$ to explain why they make a particular decision. According to  [wei_chain--thought_2022], even though the thoughts $R$ are not executed in thissystems, they help LLMs make more reasonable and feasible decisions. An example of the thought $R$ is shown below:
@@
The generated LMPs ($P$) are then sent back from the cloud to the vehicle’s electronic control unit (ECU), where they are executed. Additionally, two kinds of safety checks are set for the generated LMPs. The first check involves verifying the format of the generated code. If the code does not conform to a valid format, this system will not execute or respond to the code. The second safety check is parameter verification, which assesses the feasibility and safety of the parameters within the context of the given policy, preventing the execution of potentially risky policies. For example, if the generated code instructs our agent to move to the left lane while the ego agent is already in the leftmost lane, this system will identify this abnormality and prevent the execution.
@@ V-A5 Memory Module and Personalization
Given the adaptive nature of LLMs, if users respond differently to similar commands, the LLMs will prioritize the most recent response as a reference point for their current decision-making process. When a command from the users is issued, the LLMs access the memory module and take stored information ($H$) as part of the input prompts for the decision-making process. Additionally, each user has their own profile in the memory module, ensuring that this framework can deliver personalized driving strategies for everyone.
@@ V-B Autonomous Driving System
As shown in Fig. 6, an autonomous vehicle is used to conduct real-world experiments, which is a drive-by-wire enabled 2019 Lexus RX450h with perception sensors, localization module, and communication module. We deploy the open-source autonomous driving software Auoware.AI [autoware] with ROS Melodic in Ubuntu 18.04. We employ 3D-NDT [3Dndt] for mapping and localization, and pure pursuit [pure_pursuit] is utilized for trajectory following. Network communication is enabled by the Cradlepoint IBR900 Series Router, which is equipped with an AT&T sim card with a 4G-LTE cellular connection.
@@ V-C1 Scenarios Setup and Participants
Our experimental trials include three different scenarios: highway, intersection, and parking lot.  [^2] The experiments conducted in this study satisfy all local traffic guidelines and guarantee the safety of the participants. A human always sits in the driver’s seat of the autonomous vehicle to monitor its status and get ready to take over. The field experiments for the highway and intersection scenarios are conducted at a proving ground in Columbus, IN, USA, to validate the efficacy of the proposed Talk2Drive framework. The highway scenario involves a three-lane highway, while the intersection contains a two-way junction. Additionally, the parking lot scenario is evaluated at the North Stadium Parking Lot in West Lafayette, IN, USA. The overview visualization and statistics of the test tracks are shown in Fig. 7, while the visualization for the experiments is shown in Fig. 8. In this study, a diverse range of drivers are recruited, including seven individuals of different genders (61.4% male, 28.6% female), ages (mean=26.71, std=4.11), and driving experiences(mean=6.79, std=5.08). All participants have valid driving licenses.
@@ V-C2 Input Instructions
Given that conventionally indirect strategies primarily modify the politeness level of command of direct strategies, this section instead divides the non-conventionally indirect requests into two categories based on the strength of the hints. We define levels of directness in Tab. VII as the way to classify varying degrees of implicitness in the commands. To gather data, the test drivers are requested to generate commands based on their normal speaking preferences and subsequently categorize their commands into defined levels. Examples of the commands generated are also presented in Tab. VII.
@@ V-C3 Evaluation Metrics
where $\tau_{c}$ is the critical threshold. While 2.0 s is a common simulation margin, a 1.5 s threshold was selected for this real-world test based on existing research results on TTC value [van1994time]. For other metrics like speed variance, acceleration, and jerk, the thresholds are related to individual perception and would vary in different driving scenarios. Therefore, the score to quantify the corresponding performance is defined, with a higher score indicating better performance relative to the baseline value. For example, the sub-score for mean absolute jerk $|\bar{J}|$ can be denoted as:
@@
where $\gamma$ is the sensitivity factor, which is set empirically. The sub-scores for mean absolute acceleration and speed variance are obtained in the same way.
LLMs’ Latency for Time Efficiency: Latency $Lcy$ measures the response time of the LLM, which is important for time-sensitive vehicle applications. To measure latency, the time difference between the moment a command is sent to the LLMs and when the LLMs return the LMPs is calculated.
@@ V-D1 The Validation of Driving Performance of Talk2Drive
To validate the driving performance of the Talk2Drive system, comprehensive experiments are conducted to compare the driving performances between the baseline and the Talk2Drive through the metrics defined in Sec. V-C3. As discussed earlier, the baseline values are from the average data of human drivers. The average driving performance for each scenario within each command directness is shown in Tab. VIII. Note that for each expected driving behavior, the driving pattern is distinct. Therefore, a baseline is created for every driving behavior separately.
The driving score is the weighted sum of all the safety and comfort metrics, reflecting the overall driving performance. The driving scores are calculated and recorded in Tab. VIII, with green texts highlighting scores higher than the baseline of the same driving behavior and red texts showing the lower scores. As shown in Tab. VIII, regardless of the command’s directness or the human’s behavior, nearly all driving scores are better than the baseline, showing that this framework consistently provides relatively safe and comfortable driving experiences for humans in all scenarios.
@@
The latency of the internet remains mostly consistent (around 300 ms) under the same network conditions. The empirical results from the tests reveal that the latency for the most responsive LLM falls between approximately 1.2 to 1.8 seconds. While the observed 1.2 s to 1.8 s latency is acceptable for high-level, non-urgent decisions (e.g., responding to “drive more conservatively”), it is necessary to emphasize that this latency makes the cloud-based LLM entirely unsuitable for any safety-critical, time-sensitive maneuvers (e.g., emergency braking or collision avoidance.) This finding suggests considerable potential for integration into certain real-time applications where such response times are within acceptable limits. Specifically, LLMs demonstrate strong suitability for making high-level decisions or modifying driving behaviors in response to human commands. These scenarios require timely responses but do not typically demand extremely low latency.